how much each of 20% and 80% alcohol should be added to get 5 gallons of a 50% alcohol mixture
Answer by lwsshak3(11628)

how much each of 20% and 80% alcohol should be added to get 5 gallons of a 50% alcohol mixture
**
let x=amt of 20% alcohol to add
5-x=amt of 80% alcohol to add
..
20%x+80%(5-x)=50%*5
.2x+4-.8x=2.5
.6x=1.5
x=2.5
5-x=2.5
..
amt of 20% alcohol to add=2.5 gal
amt of 80% alcohol to add=2.5 gal